. NEED ANSWER ASAP / ANSWER NEVER USED BEFORE, COMPLETELY NEW ANSWER PLEASE

The Wendt Corporation reported $35 million of taxable income. Its federal tax rate was 21% (ignore any possible state corporate taxes).

  1. What is the company's federal income tax bill for the year? Enter your answer in dollars. For example, an answer of $1.2 million should be entered as 1,200,000. Round your answer to the nearest dollar.

    $  

  2. Assume the firm receives an additional $2 million of interest income from some bonds it owns. What is the additional tax on this interest income? Enter your answer in dollars. For example, an answer of $1.2 million should be entered as 1,200,000. Round your answer to the nearest dollar.

    $  

  3. Now assume that Wendt does not receive the interest income but does receive an additional $2 million as dividends on some stock it owns. Recall that 50% of dividends received are tax exempt. What is the additional tax on this dividend income? Enter your answer in dollars. For example, an answer of $1.2 million should be entered as 1,200,000. Round your answer to the nearest dollar.

    $  
Answer #1

Answer (a):

Company's federal income tax bill for the year = $7,350,000

Working:

Income tax bill = Taxable income * Tax rate = 35000000 * 21% = $7,350,000

Answer (b):

Additional tax on this interest income = $420,000

Working:

Additional Income tax on interest = Interest income * Tax rate = 2000000 * 21% = $420,000

Answer (c):

Additional tax on this dividend income = $210,000

Working:

Dividend Income = $2,000,000

Less: DRD = ($1,000,000)

----------------------

Taxable dividend = $1,000,000

Additional tax on this dividend income = Taxable dividend * Tax rate = 1000000 * 21% = $210,000